# Addressing Polarization and Unfairness in Performative Prediction

### v1

@@ Abstract
When machine learning (ML) models are used in applications that involve humans (e.g., online recommendation, school admission, hiring, lending), the model itself may trigger changes in the distribution of targeted data it aims to predict. Performative prediction (PP) is a framework that explicitly considers such model-dependent distribution shifts when learning ML models. While significant efforts have been devoted to finding performative stable (PS) solutions in PP for system robustness, their societal implications are less explored and it is unclear whether PS solutions are aligned with social norms such as fairness. In this paper, we set out to examine the fairness property of PS solutions in performative prediction. We first show that PS solutions can incur severe polarization effects and group-wise loss disparity. Although existing fairness mechanisms commonly used in literature can help mitigate unfairness, they may fail and disrupt the stability under model-dependent distribution shifts. We thus propose novel fairness intervention mechanisms that can simultaneously achieve both stability and fairness in PP settings. Both theoretical analysis and experiments are provided to validate the proposed method.
@@ 1 Introduction
Modern supervised learning excels in many domains, especially in pattern recognition where the target data distribution is static and independent of the deployed model. However, in the presence of model-dependent distribution shifts (i.e., the model itself triggers the changes in the distribution of targeted data), most conventional training approaches will fail to find solutions that are stable and robust to such shifts. In fact, model-dependent distribution shifts are practical and commonly exist in real-world applications. Examples include strategic individuals manipulating their data (in school admission, hiring, or lending) to game the ML system into making favorable predictions [1], consumers changing their retention and participation choices (in digital platforms) based on their perception toward the ML model they are subject to [2, 3], etc.
@@
To make predictions in the presence of model-dependent distribution shifts, Perdomo et al. [4] proposed performative prediction (PP), a framework that explicitly considers the target data distribution $\mathcal{D}(\boldsymbol{\theta})$ as a function of the ML model parameter $\boldsymbol{\theta}$ to be optimized. While PP captures the impacts of the ML model on target data, the distribution $\mathcal{D}(\boldsymbol{\theta})$ is solely determined by the model regardless of the original data distribution. A subsequent study [5] extended PP and proposed a more generalized state-dependent performative prediction (SDPP) framework, which considers the impacts of both model and initial data distribution. Specifically, given the deployed ML model parameter $\boldsymbol{\theta}$ and initial data distribution $\mathcal{D}$, SDPP models the resulting target data distribution $\mathcal{D}^{\prime}=\text{T}(\boldsymbol{\theta};\mathcal{D})$ using some transition mapping function T. Since PP is a special case of SDPP, we focus on SDPP in this paper.
Since the target data distribution depends on the model, the learning objective of SDPP is to minimize performative risk (PR), i.e.,
@@
$\boldsymbol{\theta}^{\text{PO}}=\underset{\boldsymbol{\theta}}{\operatorname{argmin}}~{}\text{PR}(\boldsymbol{\theta})\stackrel{{\scriptstyle\text{def}}}{{=}}\mathbb{E}_{Z\sim\text{T}(\boldsymbol{\theta};\mathcal{D})}[\ell(\boldsymbol{\theta};Z)]$
@@
where $\ell(\boldsymbol{\theta};Z)$ is the loss function, $Z$ is the data sampled from the target distribution $\text{T}(\boldsymbol{\theta};\mathcal{D})$ dependant on both model parameter $\boldsymbol{\theta}$ to be optimized and the initial data distribution $\mathcal{D}$. The minimizer $\boldsymbol{\theta}^{\text{PO}}$ is named as performative optimal (PO) solution. Because the target data distribution itself is a function of variable $\boldsymbol{\theta}$ to be optimized, finding the PO solution is often challenging [4, 5]. Instead, existing works have mostly focused on finding performative stable (PS) solution $\boldsymbol{\theta}^{\text{PS}}$, which minimizes the decoupled performative risk $\text{DPR}(\boldsymbol{\theta};\boldsymbol{\theta}^{\text{PS}})$ defined as follows,
@@
where $\mathcal{D}^{\text{PS}}$ is the fixed point data distribution induced by $\boldsymbol{\theta}^{\text{PS}}$ that satisfies $\mathcal{D}^{\text{PS}}=\text{T}(\boldsymbol{\theta}^{\text{PS}};\mathcal{D}^{\text{PS}})$. Unlike performative risk $\text{PR}(\boldsymbol{\theta})$ where data distribution $\text{T}(\boldsymbol{\theta};\mathcal{D})$ depends on variable $\boldsymbol{\theta}$ to be optimized, decoupled performative risk $\text{DPR}(\boldsymbol{\theta};\boldsymbol{\theta}^{\text{PS}})$ decouples the two, i.e., data distribution is induced by $\boldsymbol{\theta}^{\text{PS}}$ while the variable to be optimized is $\boldsymbol{\theta}$. Although in general $\boldsymbol{\theta}^{\text{PS}}\neq\boldsymbol{\theta}^{\text{PO}}$, $\boldsymbol{\theta}^{\text{PS}}$ is the fixed point of Eqn. (1) and stabilizes the system: at $\boldsymbol{\theta}^{\text{PS}}$, data distribution $\mathcal{D}^{\text{PS}}$ also remain fixed. Many algorithms have been proposed in the literature to find $\boldsymbol{\theta}^{\text{PS}}$. A prime example is repeated risk minimization (RRM) [4], an iterative algorithm that finds $\boldsymbol{\theta}^{\text{PS}}$ (under certain conditions) by repeatedly updating the model $\boldsymbol{\theta}^{(t)}$ that minimizes risk on the fixed distribution $\mathcal{D}^{(t-1)}$ induced by the previous model $\boldsymbol{\theta}^{(t-1)}$, i.e.,
$\boldsymbol{\theta}^{(t)}=\underset{\boldsymbol{\theta}}{\operatorname{argmin}}~{}\mathbb{E}_{Z\sim\mathcal{D}^{(t-1)}}[\ell(\boldsymbol{\theta};Z)],~{}~{}~{}\mathcal{D}^{(t)}=\text{T}(\boldsymbol{\theta}^{(t)};\mathcal{D}^{(t-1)}).$ (2)
@@
However, the societal implications of PS solutions are less understood and it is unclear whether PS solutions can cause harm and violate social norms such as fairness.
@@
In this paper, we examine the fairness properties of PS solutions $\boldsymbol{\theta}^{\text{PS}}$. We consider scenarios where an ML model is used to make decisions about people from multiple demographic groups, and the population data distribution changes based on the ML decisions they are subject to. We find that PS solutions can 1) incur severe polarization effects: entire population $\mathcal{D}^{\text{PS}}$ is dominated by certain groups, leaving the rest marginalized and almost diminished in the sample pool; 2) be biased when deployed on $\mathcal{D}^{\text{PS}}$ and people from different groups will experience different losses. However, in many applications such as recommendation systems, school admission, and loan applications, it is critical to ensure equal quality of ML predictions and population diversity. Thus, we investigate under what conditions and by what algorithms we can simultaneously achieve stability and fairness in SDPP.
@@
Focusing on group-wise loss disparity [6, 7] and participation disparity [2, 8] fairness measures, we first explore whether existing fairness mechanisms commonly used in supervised learning can help mitigate unfairness in SDPP settings; this includes regularization methods (adding fairness violation as a penalty term to the objective function of unconstrained optimization, e.g., [9, 10]) and re-weighting methods (adjusting weights and importance of samples in learning objective, e.g., [11, 12, 13]). We show that common choices of penalty terms (e.g., group-wise loss difference) and re-weighting designs (e.g., standard distributionally robust optimization) that are effective in traditional supervised learning may fail in SDPP settings by disrupting the stability of the system. Using repeated risk minimization (RRM) shown in Eqn. (2) as an example, this means that applying such fairness mechanisms at each round of RRM can disrupt the convergence of the algorithm and $(\boldsymbol{\theta}^{(t)},\mathcal{D}^{(t)})$ may diverge to an unexpected state. We thus propose novel fairness mechanisms, which can be easily adopted and incorporated into iterative algorithms such as RRM. We theoretically show that the proposed mechanism can effectively improve fairness while maintaining the stability of the system.
@@
It is worth noting that although a few recent works also studied fairness issues under model-dependent distribution shifts, they all make rather strong assumptions about the distribution shifts and do not apply to the general SDPP framework. For example, Zezulka and Genin [14], Hu and Zhang [15], Raab et al. [8] assumed there exists a causal model that depicts how data distribution would shift based on the ML model, and these causal models need to be fully known. Raab et al. [8] studied a special type of model-dependent distribution shift where only the group proportion/participation changes based on ML models. In Appendix A, we discuss these works in more detail and provide a comprehensive literature review.
@@
2 Preliminaries: Iterative Algorithms to Find PS Solutions in SDPP
As mentioned in Section 1, the original goal of SDPP is to find the performative optimal solution $\boldsymbol{\theta}^{\text{PO}}$ that minimizes the performative risk $\text{PR}(\boldsymbol{\theta})=\mathbb{E}_{Z\sim\text{T}(\boldsymbol{\theta};\mathcal{D})}[\ell(\boldsymbol{\theta};Z)]$, the loss over the population induced by the deployed model. However, solving this optimization is often challenging because the data distribution $\text{T}(\boldsymbol{\theta};\mathcal{D})$ depends on the variable $\boldsymbol{\theta}$ to be optimized. Thus, prior studies such as [4, 5] have mostly focused on finding performative stable solution $\boldsymbol{\theta}^{\text{PS}}$, which is the fixed point of Eqn. (1) and can be found through an iterative process of data sampling and model deployment. Specifically, denote $\mathcal{L}(\boldsymbol{\theta};\mathcal{D})=\mathbb{E}_{Z\sim\mathcal{D}}[\ell(\boldsymbol{\theta};Z)]$ and let $(\boldsymbol{\theta}^{(t)},\mathcal{D}^{(t)})$ be the model parameter and data distribution at round $t$ of the iterative algorithm, then repeatedly updating the model $\boldsymbol{\theta}^{(t)}$ according to Eqn. (3) could lead $(\boldsymbol{\theta}^{(t)},\mathcal{D}^{(t)})$ converging to PS solution $(\boldsymbol{\theta}^{\text{PS}},\mathcal{D}^{\text{PS}})$ under certain conditions [4, 5].
$\boldsymbol{\theta}^{(t)}=\underset{\boldsymbol{\theta}}{\operatorname{argmin}}~{}\mathcal{L}(\boldsymbol{\theta};\mathcal{D}^{(t-1)}),~{}~{}~{}\mathcal{D}^{(t)}=\text{Tr}(\boldsymbol{\theta}^{(t)};\mathcal{D}^{(t-1)}).$ (3)
where Tr is not necessarily the same as transition mapping T that drives evolution of data. Depending on how frequently the model is deployed compared to the change of data, Tr can be defined differently based on repeated deployment schema. Common examples include:
@@ 2 Preliminaries: Iterative Algorithms to Find PS Solutions in SDPP
$\mathbf{k}$-delayed: $\text{Tr}(\boldsymbol{\theta}^{(t)};\mathcal{D}^{(t-1)})=\text{T}^{k}(\boldsymbol{\theta}^{(t)};\mathcal{D}^{(t-1)})=\underbrace{\text{T}\Big{(}\boldsymbol{\theta}^{(t)};\dots\text{T}\big{(}\boldsymbol{\theta}^{(t)};\text{T}}_{k\text{ times}}(\boldsymbol{\theta}^{(t)};\mathcal{D}^{(t-1)})\big{)}\Big{)}$
@@
where the repeated risk minimization (RRM) [4] introduced in Section 1 corresponds to conventional deployment schema. By customizing the time interval between two deployments, we can get variants including delayed RRM and $k$-delayed RRM [5]. Note that the delayed deployment schema is a special case of $k$-delayed deployment schema, where the number of repeated deployments $r$ is chosen to ensure the output distribution $\mathcal{D}^{(t)}$ is sufficiently close to the fixed point distribution when $\boldsymbol{\theta}^{(t)}$ keeps being deployed on the population $\mathcal{D}^{(t-1)}$.
Technical conditions for iterative algorithms to converge to PS solutions. As shown in [4, 5], PS solutions exist and are unique only when $\ell$ and T satisfy certain requirements. Moreover, iterative algorithms introduced in Eqn. (3) can converge to the PS solution. We introduce them below, where $\Theta$, $\mathcal{Z}$, and $\triangle(\mathcal{Z})$ denote the parameter space, sample space, and space of distributions over samples.
Definition 2.1 (Strong convexity of loss).
@@
$\ell(\boldsymbol{\theta};Z)\geq\ell(\boldsymbol{\theta}^{\prime};Z)+\langle\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{\theta};Z),\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\rangle+\frac{\gamma}{2}\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\|_{2}^{2}.$
Definition 2.2 (Joint smoothness of loss).
@@
$\|\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{\theta};Z)-\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{\theta}^{\prime};Z)\|_{2}\leq\beta\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\|_{2},~{}~{}\|\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{\theta};Z)-\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{\theta};Z^{\prime})\|_{2}\leq\beta\|Z-Z^{\prime}\|_{2}.$
Definition 2.3 (Joint sensitivity of transition).
Let $\mathcal{W}_{1}$ denote the Wasserstein-1 distance measure. The transition mapping T is $\epsilon$-jointly sensitive if for all $\boldsymbol{\theta},\boldsymbol{\theta}^{\prime}\in\Theta$ and $\mathcal{D},\mathcal{D}^{\prime}\in\triangle(\mathcal{Z})$,
$\mathcal{W}_{1}(\text{T}(\boldsymbol{\theta};\mathcal{D}),\text{T}(\boldsymbol{\theta}^{\prime};\mathcal{D}))\leq\epsilon\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\|_{2},~{}~{}\mathcal{W}_{1}(\text{T}(\boldsymbol{\theta};\mathcal{D}),\text{T}(\boldsymbol{\theta};\mathcal{D}^{\prime}))\leq\epsilon\mathcal{W}_{1}(\mathcal{D},\mathcal{D}^{\prime}).$
Lemma 2.4 (The existence of a unique PS solution [4, 5]).
The SDPP problem is guaranteed to have a unique PS solution if all of the following hold: (i) Loss $\ell(\boldsymbol{\theta};Z)$ is $\gamma$-strongly convex; (ii) Loss $\ell(\boldsymbol{\theta};Z)$ is $\beta$-joint smooth; (iii) T is $\epsilon$-joint sensitive and $\epsilon(1+\beta/\gamma)<1$. Moreover, if any of these conditions is violated, the iterative algorithms may fail to converge to the PS solution.

### v2

@@ Abstract
In many real-world applications of machine learning—such as recommendations, hiring, and lending—deployed models influence the data they are trained on, leading to feedback loops between predictions and data distribution. The performative prediction (PP) framework captures this phenomenon by modeling the data distribution as a function of the deployed model. While prior work has focused on finding performative stable (PS) solutions for robustness, their societal impacts, particularly regarding fairness, remain underexplored. We show that PS solutions can lead to severe polarization and prediction performance disparities, and that conventional fairness interventions in previous works often fail under model-dependent distribution shifts due to failing the PS criteria. To address these challenges in PP, we introduce novel fairness mechanisms that provably ensure both stability and fairness, validated by theoretical analysis and empirical results [^1] Code: https://github.com/osu-srml/FairPP.
@@ 1 Introduction
Modern supervised learning has achieved remarkable success in static environments, where the data distribution remains unaffected by model deployment. However, in many real-world applications—such as digital platforms, hiring, or lending—models influence user behavior, causing feedback loops that shift the data distribution. These model-induced shifts often render standard training methods unstable or unfair, and they are prevalent in real-world applications. Examples include strategic individuals manipulating their data (in school admission, hiring, or lending) to game the ML system into making favorable predictions (hardt_strategic_2015), consumers changing their retention and participation choices (in digital platforms) based on their perception toward the ML model they are subject to (Zhang_2019_Retention; chi2022towards).
@@
To make predictions in the presence of model-induced distribution shifts, perdomo_performative_2021 proposed performative prediction (PP), a framework that explicitly considers the target data distribution $\mathcal{D}(\boldsymbol{\theta})$ as a function of the ML model parameter $\boldsymbol{\theta}\in\Theta\subset\mathbb{R}^{d}$ to be optimized in a compact domain. While PP captures the impact of ML model on target data, the distribution $\mathcal{D}(\boldsymbol{\theta})$ is solely determined by the model regardless of the original data distribution. A subsequent study (brown2020performative) extended PP and proposed a more generalized state-dependent performative prediction (SDPP) framework, which considers the impacts of both model and initial data distribution. Specifically, given the deployed ML model parameter $\boldsymbol{\theta}$ and initial data distribution $\mathcal{D}$, SDPP models the resulting target data distribution $\mathcal{D}^{\prime}=\text{T}(\boldsymbol{\theta};\mathcal{D})$ using some transition mapping function T. Since PP is a special case of SDPP, we focus on SDPP in this paper.
When the transition map T is $1$-jointly sensitive (details are in Def. 2.3), $\text{T}(\boldsymbol{\theta};\cdot)$ is contractive and repeatedly deploying $\boldsymbol{\theta}$ will cause the induced distributions to converge to a fixed point distribution $\mathcal{D}_{\boldsymbol{\theta}}$. The learning objective of SDPP is to minimize performative risk (PR) evaluated on $\mathcal{D}_{\boldsymbol{\theta}}$, i.e.,
@@
$\boldsymbol{\theta}^{\text{PO}}=\underset{\boldsymbol{\theta}}{\operatorname{argmin}}~\text{PR}(\boldsymbol{\theta})\stackrel{{\scriptstyle\text{def}}}{{=}}\mathbb{E}_{Z\sim\mathcal{D}_{\boldsymbol{\theta}}}[\ell(\boldsymbol{\theta};Z)]~~~\text{s.t.}~~\mathcal{D}_{\boldsymbol{\theta}}=\text{T}(\boldsymbol{\theta};\mathcal{D}_{\boldsymbol{\theta}})$
@@
where $\ell(\boldsymbol{\theta};Z)$ is the loss function, $Z$ is the data sampled from the fixed point distribution $\mathcal{D}_{\boldsymbol{\theta}}$. The minimizer $\boldsymbol{\theta}^{\text{PO}}$ is named as performative optimal (PO) solution. Because the target data distribution itself is a function of variable $\boldsymbol{\theta}$ to be optimized, finding $\boldsymbol{\theta}^{\text{PO}}$ is often challenging (perdomo_performative_2021; brown2020performative). Instead, existing works have mostly focused on finding performative stable (PS) solution $\boldsymbol{\theta}^{\text{PS}}$, which minimizes the decoupled performative risk $\text{DPR}(\boldsymbol{\theta};\boldsymbol{\theta}^{\text{PS}})$ defined as follows,
@@
where $\mathcal{D}^{\text{PS}}$ is the fixed point data distribution induced by $\boldsymbol{\theta}^{\text{PS}}$ that satisfies $\mathcal{D}^{\text{PS}}=\text{T}(\boldsymbol{\theta}^{\text{PS}};\mathcal{D}^{\text{PS}})$. Unlike $\text{PR}(\boldsymbol{\theta})$ where data distribution $\text{T}(\boldsymbol{\theta};\mathcal{D})$ depends on variable $\boldsymbol{\theta}$ to be optimized, $\text{DPR}(\boldsymbol{\theta};\boldsymbol{\theta}^{\text{PS}})$ decouples the two, i.e., data distribution is induced by $\boldsymbol{\theta}^{\text{PS}}$ while the variable to be optimized is $\boldsymbol{\theta}$. Although in general $\boldsymbol{\theta}^{\text{PS}}\neq\boldsymbol{\theta}^{\text{PO}}$, $\boldsymbol{\theta}^{\text{PS}}$ is the fixed point of (1) and stabilizes the system: at $\boldsymbol{\theta}^{\text{PS}}$, data distribution $\mathcal{D}^{\text{PS}}$ also remains fixed. Many algorithms have been proposed in the literature to find $\boldsymbol{\theta}^{\text{PS}}$. A prime example is repeated risk minimization (RRM) (perdomo_performative_2021), an iterative algorithm that finds $\boldsymbol{\theta}^{\text{PS}}$ (under certain conditions) by repeatedly updating the model $\boldsymbol{\theta}^{(t)}$ that minimizes risk on the fixed distribution $\mathcal{D}^{(t-1)}$ induced by the previous model $\boldsymbol{\theta}^{(t-1)}$, i.e.,
$\boldsymbol{\theta}^{(t)}$ $=$ $\underset{\boldsymbol{\theta}}{\operatorname{argmin}}~\mathbb{E}_{Z\sim\mathcal{D}^{(t-1)}}[\ell(\boldsymbol{\theta};Z)],$
$\mathcal{D}^{(t)}$ $=$ $\text{T}(\boldsymbol{\theta}^{(t)};\mathcal{D}^{(t-1)}).$ (2)
@@
However, the societal impact of PS solutions is less understood and it is unclear whether PS solutions can cause harm and violate social norms such as fairness.
@@
In this paper, we examine the fairness properties of PS solutions. We consider scenarios where an ML model is used to make decisions about people from multiple social groups, and the population data distribution changes based on the ML model. We find that $\boldsymbol{\theta}^{\text{PS}}$ can 1) incur severe polarization effects: entire population $\mathcal{D}^{\text{PS}}$ is dominated by certain groups, leaving the rest marginalized and almost diminished in the system; 2) be biased when deployed on $\mathcal{D}^{\text{PS}}$ and people from different groups will experience different losses. Because in many important domains such as job or loan applications, it is critical to ensure equal quality of ML predictions and population diversity, we investigate under what conditions and by what algorithms we can simultaneously achieve stability and fairness in SDPP.
@@
Focusing on group-wise loss disparity (martinez2020minimax; diana2021minimax; khalili2023loss) and participation disparity (Zhang_2019_Retention; Raab_Boczar_Fazel_Liu_2024) fairness measures, we first explore whether existing fairness mechanisms commonly used in supervised learning can help mitigate unfairness in SDPP settings; this includes regularization methods (adding fairness violation as a penalty term to the objective function of unconstrained optimization, e.g., (khan2023fairness; zhang2021unified)) and re-weighting methods (adjusting weights and importance of samples in learning objective, e.g., (jung2023reweighting; duchi2018learning; duchi2023distributionally)). We show that common choices of penalty terms (e.g., group-wise loss difference) and re-weighting designs (e.g., standard distributionally robust optimization) that are effective in traditional supervised learning may fail in SDPP by disrupting the stability of the system. Using repeated risk minimization (RRM) shown in (1) as an example, this means that applying such fairness mechanisms at each round of RRM can disrupt the convergence of the iterative algorithm and $(\boldsymbol{\theta}^{(t)},\mathcal{D}^{(t)})$ may diverge to an unexpected state. We thus propose novel fairness mechanisms, which can be easily adopted and incorporated into iterative algorithms such as RRM. We theoretically show that the proposed mechanism can effectively improve fairness while maintaining the stability of the system.
@@
It is worth noting that although a few recent works also studied fairness issues under model-induced distribution shifts, they all make rather strong assumptions about the distribution shifts and do not apply to the general SDPP framework. For example, mishler2022fair pointed out the fairness issues under performative settings without providing solutions to achieve fairness and stability at the same time. zezulka2023performativity; hu2022achieving; Raab_Boczar_Fazel_Liu_2024; somerstep2024algorithmic assumed there exists a causal model that depicts how data distribution would shift based on the ML model, and these causal models need to be fully known for the fairness mechanisms to work. Raab_Boczar_Fazel_Liu_2024 studied a special type of model-induced distribution shift where only the group proportion changes. In App. A, we discuss more related works.
@@
2 Preliminaries
Iterative algorithms to find PS solutions.
As mentioned in Section 1, the original goal of SDPP is to find $\boldsymbol{\theta}^{\text{PO}}$ that minimizes $\text{PR}(\boldsymbol{\theta})=\mathbb{E}_{Z\sim\mathcal{D}_{\boldsymbol{\theta}}}[\ell(\boldsymbol{\theta};Z)]$, the loss over the population induced by the deployed model. However, solving this optimization is often challenging because the data distribution $\mathcal{D}_{\boldsymbol{\theta}}$ depends on the variable $\boldsymbol{\theta}$ being optimized. Thus, prior studies such as (perdomo_performative_2021; brown2020performative) have mostly focused on finding performative stable solution $\boldsymbol{\theta}^{\text{PS}}$, which is the fixed point of Eqn. (1) and can be found through an iterative process of data sampling and model deployment. Specifically, denote $\mathcal{L}(\boldsymbol{\theta};\mathcal{D})=\mathbb{E}_{Z\sim\mathcal{D}}[\ell(\boldsymbol{\theta};Z)]$ and let $(\boldsymbol{\theta}^{(t)},\mathcal{D}^{(t)})$ be the model parameter and data distribution at round $t$ of the iterative algorithm, then repeatedly updating the model $\boldsymbol{\theta}^{(t)}$ according to Eqn. (3) could lead $(\boldsymbol{\theta}^{(t)},\mathcal{D}^{(t)})$ converging to PS solution $(\boldsymbol{\theta}^{\text{PS}},\mathcal{D}^{\text{PS}})$ under certain conditions (perdomo_performative_2021; brown2020performative).
$\boldsymbol{\theta}^{(t)}=\underset{\boldsymbol{\theta}}{\operatorname{argmin}}~\mathcal{L}(\boldsymbol{\theta};\mathcal{D}^{(t-1)}),~~\mathcal{D}^{(t)}=\text{Tr}(\boldsymbol{\theta}^{(t)};\mathcal{D}^{(t-1)}).$ (3)
where Tr may not be the same as transition map T that drives evolution of data. Depending on how frequently the model is deployed compared to the change of data, Tr is defined differently based on repeated deployment schema. Common examples include:
@@ 2 Preliminaries
$\mathbf{k}$-delayed: $\text{Tr}(\boldsymbol{\theta}^{(t)};\mathcal{D}^{(t-1)})=\text{T}^{k}(\boldsymbol{\theta}^{(t)};\mathcal{D}^{(t-1)})$
$=$ $\underbrace{\text{T}\Big(\boldsymbol{\theta}^{(t)};\dots\text{T}\big(\boldsymbol{\theta}^{(t)};\text{T}}_{k\text{ times}}(\boldsymbol{\theta}^{(t)};\mathcal{D}^{(t-1)})\big)\Big)$
@@
where the repeated risk minimization (RRM) (perdomo_performative_2021) introduced in Section 1 corresponds to conventional deployment schema. By customizing the time interval between two deployments, we can get variants including delayed RRM and $k$-delayed RRM (brown2020performative). Note that the delayed deployment schema is a special case of $k$-delayed deployment schema, where the number of repeated deployments $r$ is chosen to ensure the output distribution $\mathcal{D}^{(t)}$ is sufficiently close to the fixed point distribution when $\boldsymbol{\theta}^{(t)}$ keeps being deployed on the population $\mathcal{D}^{(t-1)}$.
Technical conditions for iterative algorithms to converge to PS solutions.
As shown in (perdomo_performative_2021; brown2020performative), PS solutions exist and are unique only when $\ell$ and T satisfy certain conditions. Moreover, iterative algorithms introduced in Eqn. (3) can converge to the PS solution. We introduce these conditions below, where $\Theta$, $\mathcal{Z}$, and $\triangle(\mathcal{Z})$ denote the parameter space, sample space, and space of distributions over samples.
Definition 2.1 (Strong convexity of loss function).
@@
$\ell(\boldsymbol{\theta};Z)\geq\ell(\boldsymbol{\theta}^{\prime};Z)+\langle\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{\theta^{\prime}};Z),\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\rangle+\frac{\gamma}{2}\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\|_{2}^{2}.$
Definition 2.2 (Joint smoothness of loss function).
@@
$\|\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{\theta};Z)-\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{\theta}^{\prime};Z)\|_{2}$ $\leq\beta\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\|_{2}$
$\|\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{\theta};Z)-\nabla_{\boldsymbol{\theta}}\ell(\boldsymbol{\theta};Z^{\prime})\|_{2}$ $\leq\beta\|Z-Z^{\prime}\|_{2}$
Definition 2.3 (Joint sensitivity of transition map).
Let $\mathcal{W}_{1}$ denote the Wasserstein-1 distance measure. The transition map T is $\epsilon$-jointly sensitive if for all $\boldsymbol{\theta},\boldsymbol{\theta}^{\prime}\in\Theta$ and $\mathcal{D},\mathcal{D}^{\prime}\in\triangle(\mathcal{Z})$, we have
$\mathcal{W}_{1}(\text{T}(\boldsymbol{\theta};\mathcal{D}),\text{T}(\boldsymbol{\theta}^{\prime};\mathcal{D}))$ $\leq\epsilon\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\|_{2}$
$\mathcal{W}_{1}(\text{T}(\boldsymbol{\theta};\mathcal{D}),\text{T}(\boldsymbol{\theta};\mathcal{D}^{\prime}))$ $\leq\epsilon\mathcal{W}_{1}(\mathcal{D},\mathcal{D}^{\prime})$
Lemma 2.4 (Existence of a unique PS solution (brown2020performative; perdomo_performative_2021)).
SDPP problem is guaranteed to have a unique PS solution if all of the following hold: (i) $\ell(\boldsymbol{\theta};Z)$ is $\gamma$-strongly convex; (ii) $\ell(\boldsymbol{\theta};Z)$ is $\beta$-joint smooth; (iii) T is $\epsilon$-joint sensitive and $\epsilon(1+2\beta/\gamma)<1$.
Lemma 2.5 (Convergence of iterative algorithms).
If conditions (i)(ii)(iii) in Lemma 2.4 are all satisfied, iterative algorithms are guaranteed to converge to the unique solution. However, there is no convergence guarantee when $\ell(\boldsymbol{\theta};Z)$ is non-convex even if both (ii) and (iii) are satisfied.
Lemma 2.4 has been shown in Theorem 8 of brown2020performative, while we prove Lemma 2.5 based on perdomo_performative_2021 in App. D.1.